# Exact non-stationary solutions of the Euler equations in two and three dimensions

### v1

@@ 2. Ideal Hydrodynamics
By polarization, it suffices to prove that
(2.5) $\langle\mathcal{A}^{-1}u,[v,u]\rangle_{L^{2}}=0$
for any $u$ and $v$ in $\mathfrak{X}_{\mu,\text{ex}}$. We handle the two- and three-dimensional cases separately.
In two dimensions, we can write $u=\nabla^{\perp}\phi$ and $v=\nabla^{\perp}\psi$ for functions $\phi$ and $\psi$ vanishing on $\partial M$ and we have $u,v]=\nabla^{\perp}\{\phi,\psi\}$, where $\{\phi,\psi\}$ is the Poisson bracket defined by $\{\phi,\psi\}\mu=d\phi\wedge d\psi$, where $\mu$ is the Riemannian area form and $\wedge$ denotes the usual wedge product. Hence (2.5) becomes
$\langle\mathcal{A}^{-1}u,[v,u]\rangle_{L^{2}}$ $=-\int_{M}g\big(\mathcal{A}^{-1}\nabla^{\perp}\phi,\nabla^{\perp}\{\phi,\psi\}\big)\,\mu$
$=-\int_{M}g\big(\nabla^{\perp}\Delta_{0}^{-1}\phi,\nabla^{\perp}\{\phi,\psi\}\big)\,\mu$
$=-\int_{M}g\big(\nabla\Delta_{0}^{-1}\phi,\nabla\{\phi,\psi\}\big)\,\mu$
$=\int_{M}\Big(-\operatorname{div}{\big(\{\phi,\psi\}\Delta_{0}^{-1}\phi\big)}+\{\phi,\psi\}\Delta\Delta_{0}^{-1}\phi\Big)\,\mu$
$=-\int_{\partial M}\{\phi,\psi\}\Delta_{0}^{-1}\phi\,\iota_{\nu}\mu+\int_{M}\phi\{\phi,\psi\}\,\mu$
$=\int_{M}\phi\,d\phi\wedge d\psi=\tfrac{1}{2}\int_{M}d(\phi^{2}\,d\psi)$
$=\tfrac{1}{2}\int_{\partial M}\phi^{2}\,d\psi=0.$
In three dimensions we have $u,v]=-\operatorname{curl}{(u\times v)}$, where $\times$ denotes the usual cross product, so that
$\langle\mathcal{A}^{-1}u,[v,u]\rangle_{L^{2}}$ $=-\int_{M}g\big(\operatorname{curl}^{-1}u,\operatorname{curl}{(u\times v)}\big)\,\mu$
$=-\int_{M}\operatorname{div}{\Big((u\times v)\times\operatorname{curl}^{-1}u\Big)}\,\mu-\int_{M}g\big(u\times v,\operatorname{curl}{(\operatorname{curl}^{-1}u)}\big)\,\mu$
$=-\int_{\partial M}g\big((u\times v)\times\operatorname{curl}^{-1}u,\nu\big)\,\iota_{\nu}\mu-\int_{M}g(u,u\times v)\,\mu.$
Now the first term vanishes since $u$ and $v$ are both tangent to the boundary, so that $u\times v$ is parallel to the unit normal $\nu$, and the cross product vanishes; meanwhile the second term vanishes by orthogonality. ∎
As an immediate consequence, we obtain the following property of the coadjoint operator.
Corollary 2.6.
We are now ready to construct the aforementioned family of $z$ satisfying the hypothesis of Theorem 2.2.
Lemma 2.7.
$K_{X}:z\mapsto K_{X}z,\quad\mathcal{A}:z\mapsto\mathcal{A}z,\quad\operatorname{ad}_{X}:z\mapsto-[X,z]$
Proof.
From Lemma 2.1 we have the $L^{2}$-orthogonal decomposition
where, for each $k\in\mathbb{Z}_{\geq 1}$, we have $\dim(E_{k})<\infty$ and $\mathcal{A}z_{k}=\alpha_{k}z_{k}$, for all $z_{k}\in E_{k}$.
The skew-symmetry of $K_{X}$ has already been established in Corollary 2.6. For $\operatorname{ad}_{X}$ note that, for all $z\in E_{k}$, we have
$\int_{M}g\big([X,z],z\big)$ $=\int_{M}g(\nabla_{X}z,z)-g(\nabla_{z}X,z)\,\mu$
$=\tfrac{1}{2}\int_{M}g\big(X,\nabla g(z,z)\big)$
$=\tfrac{1}{2}\int_{\partial M}g(z,z)g(X,\nu)\,\iota_{\nu}\mu-\tfrac{1}{2}\int_{M}(\operatorname{div}{X})g(z,z)\,\mu,$
where $\nu$ denotes the normal vector to the boundary and $\mu$ the volume-form. Here the terms vanish because $g(\nabla_{z}X,\nu)=0$ pointwise (by definition of a Killing field), and the last two terms vanish since $X$ is divergence-free and tangent to the boundary. ∎
@@ 3. Two-Dimensional Fluids
For this section we have $\mathcal{A}=\Delta$, the Hodge Laplacian. We begin by showing, via the Bochner technique, that any surface satisfying the conditions of Theorem 2.2 must have constant curvature. We refer to Petersen [28] for the notation.
Proposition 3.1.
@@
For a Killing field $X$ we have [^2] See [28, Proposition 8.1.3.].
@@
We now present examples in the canonical compact constant curvature surfaces.
Example 3.2 (Kelvin Waves on the Flat Two-Torus).
Let $(\mathbb{T}^{2},g_{\mathrm{flat}})$ be the flat two-dimensional torus and consider the Killing field $X=\partial_{x}$ which generates horizontal translations. Since the Gaussian curvature vanishes identically, Lemma 3.1 implies that $\Delta X=0$ for any Killing field. Consequently, the only admissible value of $\lambda$ in Theorem 2.2 is $\lambda=0$. On the flat torus, the eigenfunctions of the Laplacian are the Fourier modes
$\psi_{n,m}(x,y)=e^{i(nx+my)},\qquad\Delta\psi_{n,m}=(n^{2}+m^{2})\psi_{n,m},\quad n,m\in\mathbb{Z}.$
The associated divergence-free eigenfields are given by
$z_{n,m}=\nabla^{\perp}\psi_{n,m}=\partial_{y}\psi_{n,m}\,\partial_{x}-\partial_{x}\psi_{n,m}\,\partial_{y}=i\psi_{n,m}\bigl(m\partial_{x}-n\partial_{y}\bigr)$
and we have
$\partial_{x},z_{n,m}]=\nabla^{\perp}\partial_{x}\psi_{n,m}=inz_{n,m}$
so $\alpha=n^{2}+m^{2}$ and $\zeta=n$. Hence, the corresponding complex-valued solution of the Euler equations is
$U(t,x,y)=\partial_{x}+e^{-int}z_{n,m}=\partial_{x}+ie^{i\big(n(x-t)+my\big)}\bigl(m\partial_{x}-n\partial_{y}\bigr).$
Taking the real part yields
$U(t,x,y)=\partial_{x}-m\sin\bigl(n(x-t)+my\bigr)\partial_{x}+n\sin\bigl(n(x-t)+my\bigr)\partial_{y},$
solving the Euler equations (1.1). These are the standard Kelvin waves on the flat torus.
Furthermore, the imaginary part
$Y(t,x,y)=m\cos\bigl(n(x-t)+my\bigr)\partial_{x}-n\cos\bigl(n(x-t)+my\bigr)\partial_{y},$
solves the linearized Euler equations along $U$.
Lastly, $U_{\mathfrak{R}}$ is stationary if and only if $n=0$.
On domains with boundary, additional spectral constraints arise from enforcing tangency at the boundary. We now revisit Kelvin waves in the flat setting, but on the unit disk, where the rotational Killing field replaces translation invariance and Bessel functions naturally appear.
@@
Let $(D^{2},g_{\mathrm{flat}})$ be the flat two-dimensional unit disk and consider the Killing field $X=\partial_{\theta}$ which generates rotation around the origin. Again we have $\Delta\partial_{\theta}=0$. On the flat unit disk, the eigenfunctions of the Laplacian are given by
@@
where $J_{n}$ is the $n$th Bessel function and $\beta_{n,m}$ its $m$th zero. The associated divergence-free eigenfields are given by
$z_{n,m}=\nabla^{\perp}\psi_{n,m}=\frac{1}{r}\partial_{\theta}\psi_{n,m}\,\partial_{r}-\partial_{r}\psi_{n,m}\,\partial_{\theta}=\psi_{n,m}\left(\frac{in}{r}\partial_{r}-\beta_{n,m}\frac{J_{n}^{\prime}(\beta_{n,m}r)}{J_{n}(\beta_{n,m}r)}\partial_{\theta}\right)$
and
$\partial_{\theta},z_{n,m}]=\nabla^{\perp}\partial_{\theta}\psi_{n,m}=inz_{n,m}$
so $\alpha=\beta_{n,m}^{2}$ and $\zeta=n$.
Hence, the corresponding complex-valued solution of the Euler equations is
$U(t,r,\theta)=\partial_{\theta}+e^{-int}z_{n,m}=\partial_{\theta}+J_{n}(\beta_{n,m}r)e^{in(\theta-t)}\left(\frac{in}{r}\partial_{r}-\beta_{n,m}\frac{J_{n}^{\prime}(\beta_{n,m}r)}{J_{n}(\beta_{n,m}r)}\partial_{\theta}\right).$
Taking the real part yields
$U(t,r,\theta)=\partial_{\theta}-\beta_{n,m}J_{n}^{\prime}(\beta_{n,m}r)\cos(n(\theta-t))\partial_{\theta}-\frac{n}{r}J_{n}(\beta_{n,m}r)\sin(n(\theta-t))\partial_{r},$
which solves the Euler equations (1.1). These are precisely the classical Kelvin waves on the flat unit disk.
Furthermore, the imaginary part
$Y(t,r,\theta)=-\beta_{n,m}J_{n}^{\prime}(\beta_{n,m}r)\sin(n(\theta-t))\partial_{\theta}+\frac{n}{r}J_{n}(\beta_{n,m}r)\cos(n(\theta-t))\partial_{r},$
solves the linearized Euler equations along $U$.
Lastly, $U_{\mathfrak{R}}$ is stationary if and only if $n=0$.
We now turn to the positive curvature example. On the round sphere, the presence of positive curvature alters the action of the Laplacian on Killing fields, leading to nontrivial coadjoint eigenvalues.
@@
Let $(S^{2},g_{\mathrm{round}})$ be the round two-sphere and consider the Killing field $X=\partial_{\theta}$ which generates equatorial rotation. Here we have $\Delta\partial_{\theta}=2\partial_{\theta}$. The spherical harmonics are given by
@@
where $\phi\in[-\frac{\pi}{2},\frac{\pi}{2}]$ is the latitudinal parameter and $P_{n,m}$ is the associated Legendre polynomial with the standard normalization. For $m\geq 1$ the divergence-free eigenfields are then given by
$z_{n,m}=\nabla^{\perp}\psi_{n,m}=\frac{1}{\sin\phi}\left(\partial_{\phi}\psi_{n,m}\,\partial_{\theta}-\partial_{\theta}\psi_{n,m}\,\partial_{\phi}\right)=\psi_{n,m}\left(-\frac{P_{n,m}^{\prime}(\cos\phi)}{P_{n,m}(\cos\phi)}\partial_{\theta}-\frac{in}{\sin\phi}\partial_{\phi}\right).$
These are eigenfields of our adjoint
$\partial_{\theta},z_{n,m}]=\nabla^{\perp}\partial_{\theta}\psi_{n,m}=inz_{n,m}$
and, from (2.2), our coadjoint operator
$K_{\partial_{\theta}}z_{n,m}=\Delta^{-1}[z_{n,m},\Delta\partial_{\theta}]=-i\frac{2n}{m(m+1)}z_{n,m}$
so, in contrast to the flat case, we have non-zero $\lambda=\frac{2n}{m(m+1)}$, along with $\alpha=m(m+1)$ and $\zeta=n$.
Hence, the corresponding complex-valued solution of the Euler equations is
$U(t,\theta,\phi)=\partial_{\theta}+e^{in\left(\theta+\left(\frac{2}{m(m+1)}-1\right)t\right)}P_{n,m}(\cos\phi)\left(-\frac{P_{n,m}^{\prime}(\cos\phi)}{P_{n,m}(\cos\phi)}\partial_{\theta}-\frac{in}{\sin\phi}\partial_{\phi}\right).$
Taking the real part yields
$U(t,\theta,\phi)=\partial_{\theta}-\cos\left(n\left(\theta+\left(\tfrac{2}{m(m+1)}-1\right)t\right)\right)P_{n,m}^{\prime}(\cos\phi)\partial_{\theta}+\frac{n}{\sin\phi}\sin\left(n\left(\theta+\left(\tfrac{2}{m(m+1)}-1\right)t\right)\right)P_{n,m}(\cos\phi)\partial_{\phi},$
which solves the Euler equations (1.1). These are precisely the Rossby-Haurwitz waves on the round two-sphere.
Furthermore, the imaginary part
$Y(t,\theta,\phi)=-\sin\left(n\left(\theta+\left(\tfrac{2}{m(m+1)}-1\right)t\right)\right)P_{n,m}^{\prime}(\cos\phi)\partial_{\theta}-\frac{n}{\sin\phi}\cos\left(n\left(\theta+\left(\tfrac{2}{m(m+1)}-1\right)t\right)\right)P_{n,m}(\cos\phi)\partial_{\phi},$
solves the linearized Euler equations along $U$.
Lastly, $U_{\mathfrak{R}}$ is stationary if and only if $\frac{2n}{m(m+1)}=n$, that is, $n=0$ or $m=1$.
Finally, we consider the negatively curved analogue of the flat disk. In contrast to both the flat and spherical cases, negative curvature reverses the sign of the Laplacian on rotational Killing fields, yielding a distinct family of Kelvin-type waves on compact domains in hyperbolic space.
@@
$\begin{split}\psi_{n,m}(r,\theta)&=Q^{\lvert n\rvert}_{-\tfrac{1}{2}+i\beta_{n,m}}(\cosh r)e^{in\theta},\quad n\in\mathbb{Z},m\in\mathbb{Z}_{\geq 1},\\ \Delta\psi_{n,m}&=\alpha_{n,m}\psi_{n,m},\quad\text{with }\alpha_{n,m}=-\left(\tfrac{1}{4}+\beta_{n,m}^{2}\right)\end{split}$
where $Q^{\lvert n\rvert}_{-\tfrac{1}{2}+i\beta_{n,m}}$ is the associated Legendre function with $Q^{\lvert n\rvert}_{-\tfrac{1}{2}+i\beta_{n,m}}(\cosh 1)=0$ and the branch cut chosen so that the function is real on $1,\infty)$. For $m\geq 1$ the divergence-free eigenfields are then given by
$z_{n,m}=\nabla^{\perp}\psi_{n,m}=\psi_{n,m}\left(\frac{in}{\sinh r}\partial_{r}-\sinh^{2}r\frac{{Q^{\lvert n\rvert}_{-\tfrac{1}{2}+i\beta_{n,m}}}^{\prime}(\cosh r)}{Q^{\lvert n\rvert}_{-\tfrac{1}{2}+i\beta_{n,m}}(\cosh r)}\partial_{\theta}\right).$
These are eigenfields of our adjoint
$\partial_{\theta},z_{n,m}]=\nabla^{\perp}\partial_{\theta}\psi_{n,m}=inz_{n,m}$
and, from (2.2), our coadjoint operator
$K_{\partial_{\theta}}z_{n,m}=\Delta^{-1}[z,\Delta\partial_{\theta}]=-i\frac{2n}{\tfrac{1}{4}+\beta_{n,m}^{2}}z_{n,m}$
so again, in contrast to the flat case, we have non-zero $\lambda_{n,m}=\frac{2n}{\tfrac{1}{4}+\beta_{n,m}^{2}}$, along with $\alpha_{n,m}=-\left(\tfrac{1}{4}+\beta_{n,m}^{2}\right)$ and $\zeta=n$.
Hence, the corresponding complex-valued solution of the Euler equations is
$U(t,r,\theta)=\partial_{\theta}+e^{in\left(\theta+\left(\tfrac{2}{\alpha_{n,m}}-1\right)t\right)}Q^{\lvert n\rvert}_{-\tfrac{1}{2}+i\beta_{n,m}}(\cosh r)\left(\frac{in}{\sinh r}\partial_{r}-\sinh^{2}r\frac{{Q^{\lvert n\rvert}_{-\tfrac{1}{2}+i\beta_{n,m}}}^{\prime}(\cosh r)}{Q^{\lvert n\rvert}_{-\tfrac{1}{2}+i\beta_{n,m}}(\cosh r)}\partial_{\theta}\right).$
Taking the real part yields
$\begin{split}U(t,r,\theta)&=\partial_{\theta}-\sinh^{2}r\cos\left(n\left(\theta+\left(\tfrac{2}{\alpha_{n,m}}-1\right)t\right)\right){Q^{\lvert n\rvert}_{-\tfrac{1}{2}+i\beta_{n,m}}}^{\prime}(\cosh r)\partial_{\theta}\\ &\qquad\qquad-\frac{n}{\sinh r}\sin\left(n\left(\theta+\left(\tfrac{2}{\alpha_{n,m}}-1\right)t\right)\right)Q^{\lvert n\rvert}_{-\tfrac{1}{2}+i\beta_{n,m}}(\cosh r)\partial_{r},\end{split}$
which solves the Euler equations (1.1). These are the hyperbolic analogue of Kelvin waves on the flat unit disk.
Furthermore, the imaginary part
$\begin{split}Y(t,r,\theta)&=-\sinh^{2}r\sin\left(n\left(\theta+\left(\tfrac{2}{\alpha_{n,m}}-1\right)t\right)\right){Q^{\lvert n\rvert}_{-\tfrac{1}{2}+i\beta_{n,m}}}^{\prime}(\cosh r)\partial_{\theta}\\ &\qquad\qquad+\frac{n}{\sinh r}\cos\left(n\left(\theta+\left(\tfrac{2}{\alpha_{n,m}}-1\right)t\right)\right)Q^{\lvert n\rvert}_{-\tfrac{1}{2}+i\beta_{n,m}}(\cosh r)\partial_{r},\end{split}$
solves the linearized Euler equations along $U$.

### v2

@@ 2. Ideal Hydrodynamics
From Proposition 2.6, we have an entire basis of complex fields $\mathfrak{q}\in\mathbb{C}\otimes\mathfrak{X}_{\mu,\text{ex}}$ satisfying the assumptions of Theorem 2.2 with $u_{0}=X$. Decomposing each as $\mathfrak{q}=v+iw$ with $v,w\in\mathfrak{X}_{\mu,\text{ex}}$ and recalling the definition of the operator $K_{X}$ from (2.4) completes the proof.
Remark 2.8.
Note that locally we can construct polar coordinates such that $X=\partial_{\theta}$ which in turn induces a traveling wave type structure on our solutions in the corresponding chart; that is, the dependence on $\theta$ is in terms of $(\theta-t)$. It is conceivable however that no such global coordinate $\theta$ exists, depending on whether the orbits of $X$ are closed and of the same length.
@@ 3. Two-Dimensional Fluids
For this section we have $\mathcal{A}=\Delta$, the positive-definite Hodge Laplacian on vector fields, cf. (2.2). We begin by showing, via the Bochner technique, that any surface satisfying the conditions of Theorem 2.2 must have constant curvature. The proof can also be carried out in local coordinates, but the technique of Bochner yields a global result more readily. We refer to Petersen [31] for the notation.
Theorem 3.1.
@@
For a Killing field $X$ we have [^3] See [31, Proposition 8.1.3.].
@@
Note that when the fluid domain is a compact surface of constant sectional curvature $\kappa$, we can express the metric locally in polar coordinates
$g=dr^{2}+s_{\kappa}(r)d\theta^{2}.$
where
(3.1) $s_{\kappa}(r)=\begin{cases}\sin{r}&\kappa=1,\\ r&\kappa=0,\\ \sinh{r}&\kappa=-1,\end{cases}$
and $X=\partial_{\theta}$ is Killing with $\Delta X=2\kappa X$.
The following theorem outlines the general scheme for applying Theorem (2.2) in this setting.
Theorem 3.2.
Let $(M,g)$ be a surface of constant sectional curvature $\kappa$. The eigenfunctions of the Laplacian are of the form
$\psi_{n,m}(r,\theta)=e^{in\theta}F_{n,m}(r),\qquad\Delta\psi_{n,m}=\alpha_{n,m}\psi_{n,m},\quad n\in\mathbb{Z},m\in\mathbb{Z}_{\geq 0}$
and, defining $\varpi_{n,m}=n\left(\frac{2\kappa}{\alpha_{n,m}}-1\right)$, the vector field
(3.2) $U_{\mathfrak{R}}(t,r,\theta)=\partial_{\theta}+\frac{F_{n,m}^{\prime}(r)}{s_{\kappa}(r)}\cos\left(n\theta+\varpi_{n,m}t\right)\partial_{\theta}+\frac{nF_{n,m}(r)}{s_{\kappa}(r)}\sin\left(n\theta+\varpi_{n,m}t\right)\partial_{r}$
solves the Euler equations (1.1).
Furthermore $U_{\mathfrak{R}}$ is stationary if and only if $\varpi_{n,m}=0$.
Proof.
Taking the skew-gradient of the Laplace-eigenfunctions, we acquire divergence-free fields
${\mathfrak{q}}_{n,m}=\nabla^{\perp}\psi_{n,m}=\frac{1}{s_{\kappa}(r)}\big(-\partial_{\theta}\psi_{n,m}\partial_{r}+\partial_{r}\psi_{n,m}\partial_{\theta}\big)=\frac{e^{in\theta}}{s_{\kappa}(r)}\big(-inF_{n,m}(r)\partial_{r}+F_{n,m}^{\prime}(r)\partial_{\theta}\big)$
satisfying
$\Delta\mathfrak{q}_{n,m}=\alpha_{n,m}\mathfrak{q}_{n,m}.$
These are eigenfields of our adjoint
$\partial_{\theta},{\mathfrak{q}}_{n,m}]=\nabla^{\perp}\partial_{\theta}\psi_{n,m}=in{\mathfrak{q}}_{n,m}$
and, from (2.4), our coadjoint operator
$K_{\partial_{\theta}}{\mathfrak{q}}_{n,m}=\Delta^{-1}[{\mathfrak{q}}_{n,m},\Delta\partial_{\theta}]=-i\frac{2\kappa n}{\alpha_{n,m}}{\mathfrak{q}}_{n,m}.$
Hence, in the context of Theorem 2.2, we have $\lambda_{n,m}=\frac{2\kappa n}{\alpha_{n,m}}$ and $\zeta_{n,m}=n$.
Letting $\varpi_{n,m}=\lambda_{n,m}-\zeta_{n,m}=n\left(\frac{2\kappa}{\alpha_{n,m}}-1\right)$, the corresponding complex-valued solution of (1.1) is given by
$U_{\mathfrak{C}}(t)=\partial_{\theta}+e^{i\varpi_{n,m}t}\mathfrak{q}_{n,m}.$
Taking the real part yields (3.2). ∎
We now consider some canonical compact constant curvature surfaces.
@@
Let $(D^{2},g_{\mathrm{flat}})$ be the flat two-dimensional unit disk and consider the Killing field $X=\partial_{\theta}$ which generates rotation around the origin. Here we have $\kappa=0$, and hence $\Delta\partial_{\theta}=0$. Consequently, the only admissible value of $\lambda$ in Theorem 2.2 is $\lambda=0$. Hence, the solutions we generate in this setting are trivial in the sense outlined in Remark 2.4.
On the flat unit disk, $s_{\kappa}(r)=r$ and the eigenfunctions of the Laplacian are given by
@@
where $J_{n}$ is the $n^{\text{th}}$ Bessel function and $\beta_{n,m}$ its $m^{\text{th}}$ zero. Letting $\varpi_{n,m}=-n$, the corresponding solutions (3.2) are precisely the classical Kelvin waves on the flat unit disk.
In Section 4 we present a three-dimensional analogue of the above where the helical symmetry creates a twisting effect which generates nontrivial flows, cf. Example 4.13.
On the round sphere, the presence of positive curvature alters the action of the Laplacian on Killing fields, leading to nontrivial eigenvalues $\lambda$.
@@
Let $(S^{2},g_{\mathrm{round}})$ be the round two-sphere and consider the Killing field $X=\partial_{\theta}$ which generates equatorial rotation. Here we have $\Delta\partial_{\theta}=2\partial_{\theta}$ and, replacing $r$ with the more natural $\phi\in[0,\pi]$ denoting the latitudinal parameter, $s_{\kappa}(\phi)=\sin(\phi)$. The spherical harmonics are given by
@@
where $P_{n,m}$ is the associated Legendre polynomial with the standard normalization. Letting $\varpi_{n,m}=n\left(\frac{2}{m(m+1)}-1\right)$, the corresponding solutions (3.2) are precisely the Rossby-Haurwitz waves on the round two-sphere. Furthermore, $U_{\mathfrak{R}}$ is stationary if and only if $\frac{2n}{m(m+1)}=n$, that is, $n=0$ or $m=1$.
The simplest non-stationary example is then $n=1$ and $m=2$. Here, we get that $P_{1,2}(\cos{\phi})$ is a multiple of $\sin{\phi}\cos{\phi}$ and $\varpi_{1,2}=-\tfrac{2}{3}$. The corresponding solution from (3.2) is given by
(3.3) $U_{\mathfrak{R}}(t,\phi,\theta)=\partial_{\theta}+\frac{\cos(2\phi)}{\sin(\phi)}\cos\left(\theta-\frac{2}{3}t\right)\partial_{\theta}+\cos(\phi)\sin\left(\theta-\frac{2}{3}t\right)\partial_{\phi}$
which can be expressed in Cartesian coordinates as the restriction to the sphere of
(3.4) $U_{\mathfrak{R}}(t)=e_{1}-\left(xe_{1}+ze_{2}\right)\cos\left(\frac{2t}{3}\right)-\left(ye_{1}+ze_{3}\right)\sin\left(\frac{2t}{3}\right)$
where
$e_{1}=-y\partial_{x}+x\partial_{y},\quad e_{2}=-z\partial_{y}+y\partial_{z},\quad e_{3}=-x\partial_{z}+z\partial_{x}.$
In Section 4 we present a three-dimensional analogue of these solutions on $S^{3}$, cf. (4.4).
Finally, we consider the negative curvature case. Note that by Bochner’s theorem (cf. Petersen [31, Chapter 7]), a closed manifold with negative Ricci curvature admits no nontrivial Killing fields. Hence, our examples in this setting are restricted to surfaces with boundary. We will focus on the hyperbolic analogue of the flat disk. In contrast to both the flat and spherical cases, negative curvature reverses the sign of the Laplacian on rotational Killing fields, yielding a distinct family of Kelvin-type waves on compact domains in hyperbolic space.
@@
$\psi_{n,m}(r,\theta)=Q_{n,m}(\cosh{r})e^{in\theta},\qquad\Delta\psi_{n,m}=\alpha_{n,m}\psi_{n,m},\quad\text{with }\alpha_{n,m}=\tfrac{1}{4}+\beta_{n,m}^{2},\quad n\in\mathbb{Z},m\in\mathbb{Z}_{\geq 1},$
where $Q_{n,m}:=Q^{\lvert n\rvert}_{-\tfrac{1}{2}+i\beta_{n,m}}$ is the associated Legendre function [^4] For details on these eigenfunctions, see [37]. with $\beta_{n,m}$ such that $Q_{n,m}(\cosh{1})=0$ and the branch cut chosen so that the function is real on $1,\infty)$. Letting $\varpi_{n,m}=n\left(\frac{2}{\alpha_{n,m}}-1\right)$, the corresponding solutions (3.2) are a hyperbolic analogue of Rossby-Haurwitz waves.